OpenStudy (tester97):

Garrett throws a dart at a circular dart board. The dart board has a radius of 18 inches, and the bull’s eye in the center of the dart board has a radius of 4 inches. What is the probability that a dart thrown at random within the dartboard will hit the bull’s eye? Round your answer to the nearest tenth, if necessary.

hartnn (hartnn):

Find the area of the board (this will be your sample space) and area covered by the bull's eye (this will be numerator for calculating probability)

OpenStudy (tester97):

How would i do that?

hartnn (hartnn):

area of circle = \(\Large \pi r^2\)

hartnn (hartnn):

for the board, radius = 18 for the bull's eye, radius = 4

OpenStudy (tester97):

so i would do it like this? \[\frac{ 4^2\pi }{ 18^2\pi }\]

hartnn (hartnn):

yes, correct. pi gets cancelled out

hartnn (hartnn):

if you want your final answer in %, then multiply the above result by 100

OpenStudy (tester97):

so this? \[\frac{ 4 }{ 18 }\times100?\]

hartnn (hartnn):

no... did you ignore the squares ? \(\Large \frac{ 4 ^2}{ 18^2 }\times100\)

OpenStudy (tester97):

Oh yea sorry i forgot

OpenStudy (tester97):

but i got 4.93 is that right?

hartnn (hartnn):

4.938% or 4.94 % yes thats correct!

OpenStudy (tester97):

Thanks Hartnn ;)

hartnn (hartnn):

welcome ^_^